OpenStudy (anonymous):

Craig feeds his pets each night. He feeds his big dog 1/2 a can of dog food. and feeds his small dog 1/3 can of dog food. How much dog food does he need each night?

OpenStudy (anonymous):

\[\frac{1}{2}\text{ can}+\frac{1}{3}\text{ can}=\frac{5}{6}\text{ can}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome and thank you for the medal.